Translate the problem into a pair of linear equations in two variables.
Jill is 11km away from Joe; both begin to walk toward each other at the same time Jill walks at 2.5km/hr. They meat in 2 hours how fast is Joe walking.

Answer by checkley77(12844)

D=RT
11=(2.5+X)2
11=5+2X
2X=11-5
2X=6
X=6/2
X=3 KM/HR.IS JOE'S SPEED.
PROOF:
11=(2.5+3)2
11=5.5*2
11=11